OpenStudy (anonymous):

what is the value of x in the equation of 3/5=x/30

jimthompson5910 (jim_thompson5910):

Cross multiply to go from 3/5 = x/30 to 3*30 = 5*x

jimthompson5910 (jim_thompson5910):

Let me know what you get when you solve that

OpenStudy (anonymous):

ok i got 18

jimthompson5910 (jim_thompson5910):

you got it

OpenStudy (anonymous):

but on my homework it says 6 or 10 or 9 or 180 is it 180?

jimthompson5910 (jim_thompson5910):

hmm there's a zero that's missing (that's not in the original problem you wrote)

jimthompson5910 (jim_thompson5910):

check your problem again

OpenStudy (anonymous):

i did it is 3/5=x/10

jimthompson5910 (jim_thompson5910):

oh its 10 and not 30?

OpenStudy (anonymous):

thank u so much

jimthompson5910 (jim_thompson5910):

so what did you get?

OpenStudy (anonymous):

how did u get it. explain what u did

jimthompson5910 (jim_thompson5910):

3/5=x/10 becomes 3*10 = 5*x after cross multiplying

jimthompson5910 (jim_thompson5910):

tell me what you get when you solve

OpenStudy (anonymous):

i got 6

jimthompson5910 (jim_thompson5910):

that's correct